OpenStudy (anonymous):

Estimate using compatible numbers. 49% of 609 "PLEASE HELP IT'S TO PASS MY FINAL EXAM!" Thanks Answer(S) 240 300 350 410

OpenStudy (anonymous):

\[=49/100 \times 609\]\[\rightarrow 50/100 \times 610 = 1/2 \times 610= 305\] this way?